OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

Could you please help me @iGreen

OpenStudy (igreen):

Parallel lines have the same slope..so we can eliminate two options already..which options are those?

OpenStudy (igreen):

Do you know what the slope of y = -3x + 12 is? It's in the form of y = mx + b, where m = slope..so you can tell what the slope is just by looking at it.

OpenStudy (igreen):

@yoyogators

OpenStudy (anonymous):

So the slope is -3?

OpenStudy (igreen):

Yep. So parallel lines have the same slope, so you can get rid of two options..the one's that don't have a slope of -3.

OpenStudy (igreen):

We can eliminate Options A and C because they have a slope of 1/3.

OpenStudy (igreen):

That leaves us with B and D. y = -3x + 3 and y = -3x + 7 We can plug in (-1, 6) into each of these and solve and see which one makes it true. y = -3x + 3 6 = -3(-1) + 3 Multiply: 6 = 3 + 3 Can you add 3 + 3? @yoyogators

OpenStudy (anonymous):

So it would be B?

OpenStudy (igreen):

Yep, because the equation matches the point, and it has the same slope.

OpenStudy (anonymous):

Okay i see Thanks so much! :)